OpenStudy (anonymous):

f(x) = 1/(x^5-4 x). Then f' =

OpenStudy (anonymous):

f'= derivative of first times second minus derivative of second times first divided by the whole thing squared.

OpenStudy (xishem):

\[f(x)=\frac{1}{x^5-4x}\]We need to use the quotient rule...\[f'(x)=\frac{(x^5-4x)(0)-(1)(5x^4-4)}{(x^5-4x)^2}=\frac{5x^4-4}{(x^5-4x)^2}\]

OpenStudy (anonymous):

what else do i need to do after that?

OpenStudy (xishem):

That's it. That's the derivative of f(x).

Directrix (directrix):

or f ' (x) =[ (4 - 5 x^4)] / [x^2 (x^4 -4) ^2]

OpenStudy (xishem):

Directrix is right. I forgot to distribute through the -1. My answer was wrong.